Find the area.
leg is 12
leg is x+4
hypotenuse is 2x

Answer by mananth(16949)

one leg 12
another leg x+4
Hypotenuse = 2x
Area of a right triangle = = 1/2 * one leg * second leg
Area = 1/2 * 12*(x+4)
Area = 6x +24
..
..
(2x)^2= 12^2 + (x+4)^2
4x^2 = 144 +x^2 +8x+16
3x^2-8x-160=0
Use quadratic equation formula for roots x1,x2
x1= -(-8) + sqrt (-8)^2-4*3*(-160) / 2*3
x1= 8 + sqrt 64+1920 / 6
x1= 8.75
..
x2 =8 - sqrt 64+1920 / 6
x2= -6.09 ( this is not possible)
x= 8.75
Area= 6x+24
plug the value of x
6*8.75+24
= 76.5 square units